OpenStudy (anonymous):

Help? The volume of a pyramid with a rectangular base is V = 1 third times L times W times H. Find W if V = 500, L = 15, and H = 4.

jimthompson5910 (jim_thompson5910):

V = (1/3)LWH 500 = (1/3)*(15)*W*(4) 500 = 20W 500/20 = W 25 = W W = 25

OpenStudy (turingtest):

\[V={1\over3}lwh\] \[w={3V \over lh}\]plug in and solve

OpenStudy (anonymous):

Thanks guys :P